# Evaluating Automated Radiology Report Quality through Fine-Grained Phrasal Grounding of Clinical Findings

### v1

@@ 4 Results
Evaluating through proposed measure
Next, we evaluated the report quality using our lexical measure reported in Section 3 using prefix patterns restricting to core finding, anatomy (with laterality), severity. The result is shown for the 3 report generators evaluated in Table 1. From this table, we observe that the RGRG report generator has the highest lexical quality with their FFL patterns matching closely with the ground truth FFL patterns at all levels of granularity. We also notice that all methods improve in report quality when evaluated on the basis of their core finding. The Mean IOU scores were evaluated for the FFL prefixes that retained the anatomical location and are as shown in the last column of that table.
Comparison to evaluation scores:
To compare our approach to other report evaluation scores, we selected representative methods for word overlap scores (BLEU[6]), semantic textual matching (BERTscore[8]) and clinical accuracy F1-score (Radgraph[9]). The result is shown in Table 2 from which we observed that the lexical comparison scores under-estimated the accuracy of the reports due to lexical mismatch in the reported descriptions. The clinical accuracy score, as it was trained on fewer findings (14 findings), overestimated the performance by giving higher scores due to missed findings in their model. Our approach gave balanced estimates of report quality indicating 36-48% spatial overlap of their locations and 33-44% overlap in their descriptions. Finally, we observe that all reporting metrics rated the RGRG report as the best even in this evaluation.

### v2

@@ 4 Results
Evaluation through the proposed measure
Next, we evaluated the report quality using our lexical measure reported in Section 3 using prefix patterns restricting to core finding, anatomy (with laterality), severity. The result is shown for the 3 report generators evaluated in Table 1. From this table, we observe that the RGRG report generator has the highest lexical quality with their FFL patterns matching closely with the ground- truth FFL patterns at all levels of granularity. We also notice that all methods improved in report quality when evaluated on the basis of their core finding. The Mean IOU scores were evaluated for the FFL prefixes that retained the anatomical location and are as shown in the last column of that table.
Comparison with evaluation scores:
To compare our approach with other report evaluation scores, we selected representative methods for word overlap scores (BLEU[6]), semantic textual matching (BERTscore[8]) and clinical accuracy F1-score [9]. The result is shown in Table 2 from which we observed that the lexical comparison scores under-estimated the accuracy of the reports due to lexical mismatch in the reported descriptions. The clinical accuracy score, as it was trained on fewer findings (14 findings), overestimated the performance by giving higher scores due to missed findings in their model. Our approach gave balanced estimates of report quality indicating 36-48% spatial overlap of their locations and 33-44% overlap in their descriptions. Finally, we observe that all reporting metrics rated the RGRG report as the best even in this evaluation.